OpenStudy (anonymous):

In my book it says that my answer is wrong, but I am not sure how. @hartnn it says that the answer is no solution but this is what I got. |2x-3|=-1 2x-3=1 or 2x-3=1 +3 +3 +3 +3 2x=4 2x=2 /2 /2 /2 /2 x=2 or x1

OpenStudy (anonymous):

|2x-3| is either (2x-3) or -(2x-3)

OpenStudy (anonymous):

@phanta_seea I don't understnad what you are saying. It is absolute vaule.

OpenStudy (anonymous):

yes it is an absolute bracket. I believe you set up the equations like this: 2x-3=-1 or 2x-3=1 ?

hartnn (hartnn):

you forgot to verify!!

hartnn (hartnn):

put x=2 in original equation see if left side = right side or not

hartnn (hartnn):

oh, and you could have figured out the answer in one glance. |....| can NEVER be negative!! like |-4| =4 so, |2x-3| can NEVER be = -1 so infact there is NO SOLUTION

OpenStudy (anonymous):

@hartnn Ohhhh. I understand what I did! Thank you!

hartnn (hartnn):

welcome ^_^